Kim sells a small magazine full of motivational quotes to runners of all ages for $2.00 per copy. Hiring the printing press for one day, the only fixed cost is $300 an issue. The variable cost of printing each issue is $1.00 per copy. Kim is earning a profit of $500 per issue. How many copies per issue (units) is Kim selling?

What is the contribution per unit?

If Kim increases her desired profit to $1,000 per issue, how many copies will Kim need to sell?

How many units must Kim sell to break even?

David sells decorative stone garden statues, his annual Fixed Costs are $1,000 and the Variable Costs are $50 per statue. The Market Size for the garden statues in his area is 2,000 statues, and David wants to achieve a Market Share of 4%. At a price of $70 a statue, what is the unit contribution ($) per statue sold?

How many statues must David sell to break even?

If David wants to achieve a profit of $1,000, how many statues must David sell?

Poppy’s bakes and sells muffins. The company is conducting a price test to find a better price point. Presently their muffins sell for $20 per dozen. Their current volume is 1,000 dozen per month. They are considering reducing their sales price by 10% per dozen. If Poppy’s variable costs are $10 per dozen, what is their contribution per unit at current prices?

Will Poppy’s break even with fixed costs equal to $1,500?

How many dozens does Poppy’s need to sell to generate $1,000 in profit?
Answer #1

a) Kims Magazine

Details available

1) Selling price = $2

2) Fixed Cost per issue = $300

3) Variable Cost = $1

i) If Kim is making $500 per issue then the equation would be.

Profit+Variable Cost+Fixed Cost = Sale

Assuming units is x equation would be

500+1x+300 = 2x

i.e 2x-1x = 500+300

i.e x = 800 copies

ii) Contribution per unit = Sales per unit- variable cost p.u

= $2- $1

= $ 1

iii) If desire profit level is $1000, the equation would

(Fixed Costs+ Profit) / Contribution = (300+1000)/1 = 1300 Copies

2) David's Decorative Statue.

  Details given

Fixed Costs = $1000

Variable Costs = $50 per Statue

Market share required = 4% ot 2000 = 80

Sale price = $70 per statue

i) Unit Contribution = Sale price p.u - Variable Cost p.u

= $70-$50

= $ 20   

ii) Breakeven = Fixed Costs/ Contribution p.u = 1000/20 = 50 units

iii) If desire profit level is $ 1000, equation would be

(Fixed Costs+ Profit) / Contribution = (1000+1000)/20 = 100 units

3) Poppys Muffins

Details given

Current Sale price = $20 per dozen

Current sales volume = 1000 dozen

Reduction in Selling price by 10% considered.. Revised Sale price = $18

Variable Costs = $10 per dozen

i) Contribution at current prices = Current Sale price p.u - Variable Costs p.u = $20-$10 = $10

ii) Breakeven qty = Fixed Costs/ Contribution p.u

IF Fixed costs = $1500 equation would be 1500/10 = 150 dozen. Since Poppy is selling 1000 dozen p.m, Poppy is making profits

iii) If desire profit level is $ 1000, equation would be

(Fixed Costs+ Profit) / Contribution = (1500+1000)/10 = 250 dozen